What are the decimal numbers for 159, 150, 200, 113, 225, 87, 106, 81 when converted to 8-bit binary?
Answer #1


1)  159
159/2 = 79, remainder is 1
79/2 = 39, remainder is 1
39/2 = 19, remainder is 1
19/2 = 9, remainder is 1
9/2 = 4, remainder is 1
4/2 = 2, remainder is 0
2/2 = 1, remainder is 0
1/2 = 0, remainder is 1
Read from the bottom (MSB) to top (LSB) as 10011111
Answer: 10011111

2)  150
150/2 = 75, remainder is 0
75/2 = 37, remainder is 1
37/2 = 18, remainder is 1
18/2 = 9, remainder is 0
9/2 = 4, remainder is 1
4/2 = 2, remainder is 0
2/2 = 1, remainder is 0
1/2 = 0, remainder is 1
Read from the bottom (MSB) to top (LSB) as 10010110
Answer: 10010110

3)  200
Answer: 11001000

4)  113
Answer: 01110001

5)  225
Answer: 11100001

6)  87
Answer: 01010111

7)  106
Answer: 01101010

8)  81
Answer: 01010001